# thanks: Authors of equal contributionsthanks: Authors of equal contributions

### v2

@@ Introduction.
For fermionic environments, the DQME-SQ theory can be established in a similar way. For a single-level system, we consider a common bilinear form of the system-environment coupling Hamiltonian, $H_{\mbox{\tiny S}\mbox{\tiny E}}=\hat{c}^{+}\hat{F}^{-}+\hat{F}^{+}\hat{c}^{-}$, which describes the exchange of fermionic particles between the system and environment. The DD for the environment hybridization operators reads
@@
Here, $\langle\hat{f}_{k}^{\sigma}(t)\hat{f}_{k^{\prime}}^{\bar{\sigma}}(0)\rangle_{\mbox{\tiny D}}=\delta_{kk^{\prime}}\eta_{k}^{\sigma}e^{-\gamma_{k}^{\sigma}t}$ with $\sigma=\pm$ and $\bar{\sigma}=-\sigma$. Such a DD protocol completely recovers the environment correlation functions. The fermionic dissipaton Fock space is ${\cal H}_{\mbox{\tiny D}}=\otimes_{k}({\cal H}_{f^{-}_{k}}\otimes{\cal H}_{f^{+}_{k}})$, which involves both the left and right actions of dissipaton creation and annihilation operators. Thus, the second quantization of fermionic dissipatons is realized through the following mappings.
@@
where $\hat{N}_{k}\equiv\hat{b}_{k}^{+}\hat{b}_{k}^{-}$ is the fermionic dissipaton number operator. Details on the analytic derivation are provided in Ref. [Li23SM, ]. The compactness of fermionic DQME-SQ theory is consistent with that of the bosonic case.
The DQME-SQ theory is in principle exact and universally applicable to both bosonic and fermionic environments, accommodating arbitrary spectral density and any temperature. Consistent with the HEOM theory, it captures the complete non-Markovianity. More importantly, the DQME-SQ offers a notable advantage: its formulation embraces a compact time-local structure within a minimized Hilbert space. This unique feature enables easy encoding into quantum circuits, distinguishing it from HEOM. While the computational complexity remains comparable to HEOM when implemented with classical algorithms, its excellent compatibility with quantum algorithms makes DQME-SQ a preferred choice for simulating complex non-Markovian OQS.
@@
Quantum simulation of DQME-SQ.
To simulate the non-Markovian open quantum dynamics governed by the DQME-SQ using quantum circuits, the first key step is encoding the extended density tensor. For both bosonic and fermionic environments, the extended density tensor $\tilde{\rho}$ is mapped onto a pure qubit state, $\lvert\tilde{\rho}\rangle$, as
$\tilde{\rho}\longmapsto|\tilde{\rho}\rangle=\frac{1}{C}\sum_{ij{\bf n}}P_{ij}^{\bf n}\,|i\rangle\otimes|j\rangle\otimes|\bf n\rangle.$ (8)
Here, the summation is over all the possible configurations of the extended system, where $|{\bf n}\rangle=|n_{1}\rangle\otimes|n_{2}\rangle\cdots\otimes|n_{K}\rangle$ represents a dissipaton configuration. Note that for fermionic environments, $|{\bf n}\rangle$ is formed by paired dissipaton Fock states. These configurations are encoded onto the qubits using a standard binary code. The global normalization coefficient $C$ is determined by the probability conservation condition of $\sum_{i}P_{ii}^{\bf 0}=1$. The temporal evolution of $\tilde{\rho}(t)$ is then simulated by propagating the following equation of motion,
@@
where $\Lambda$ is a non-Hermitian dynamical generator that represents the right-hand side of Eq. (Construction of DQME-SQ theory.) or Eq. (Construction of DQME-SQ theory.) Li23SM .
The temporal propagation of $|\tilde{\rho}(t)\rangle$ can be realized with various quantum algorithms, including the linear combination of unitaries (LCU) Lon06825 ; Sch22023216 , quantum singular value transformation, time-dependent variational principle Sch21270503 ; Sch22023216 ; Hea21013182 ; Wan234851 , and others Lam23Arxiv2310_12539 ; Zha24054101 . Here, we utilize the LCU propagation Sch22023216 . Within a small time step $\delta t$, the propagator $e^{-i{\Lambda}\delta t}$ is approximately decomposed as $e^{-i{\Lambda}\delta t}=U_{1}(\delta t)U_{0}(\delta t)$. Here, $U_{0}(\delta t)\equiv e^{-i\Lambda_{0}\delta t}$ governed by a Hermitian generator, $\Lambda_{0}\equiv(\Lambda+\Lambda^{\dagger})/2$, is a unitary propagator, while $U_{1}(\delta t)\equiv e^{-i\Lambda_{1}\delta t}$ governed by an anti-Hermitian generator, $\Lambda_{1}\equiv(\Lambda-\Lambda^{\dagger})/2$, is non-unitary. The latter is further approximated by
@@
Shown in Fig. 2 is the designed quantum circuit of DQME-SQ comprising the target bits to store the density tensor state as well as a control bit for propagation. The input of the circuit is $|\tilde{\rho}(t)\rangle\otimes|0\rangle$. From the left to the right in the circuit, we first act a Hadamard gate (H) on the control bit, followed by the $U_{0}(\delta t)$ gate on the target bits. Subsequently, the states go through the 0-control and 1-control gates of $U^{\epsilon}_{\pm}(\delta t)$ successively. After acting an ${\rm R}_{\rm Y}(-{\pi}/{2})$ gate on the control bit, we obtain $2\epsilon|\tilde{\rho}(t+\delta t)\rangle\otimes|0\rangle+|{\rm unwanted}\rangle\otimes|1\rangle$. Finally, by performing an Z-measurement on the control bit, we project out the unwanted state and obtain the target state $|\tilde{\rho}(t+\delta t)\rangle$ with a probability depending on $\epsilon$. Once obtained, the wanted state will serve as the initial state for the next propagation step.
@@
In contrast to classical algorithms whose computational costs grow exponentially as the system size or the complexity of non-Markovianity increases Tan20020901 , the LCU algorithm exhibits highly favorable scalability – both the number of qubits and the circuit depth grow linearly with the increase of the number of dissipatons. This favorable scalability makes the DQME-SQ theory, implemented with the LCU algorithm, as a promising approach to efficiently simulate complex non-Markovian open quantum dynamics.
@@
Figure 3: (a) Simulated time evolution of $P_{1}-P_{0}$, the population difference between two states, in a spin-boson model, by using the quantum algorithm for the DQME-SQ theory, compared to low and high temperature results via the classical propagation of HEOM; (b) Simulated time evolution of the ground-state population $P_{0}$ for a single-impurity Anderson model, with inset depicting the imaginary part of the correlation function. Detailed parameters are provided in SM Li23SM .
Numerical demonstrations.

### v3

@@ III.1 Construction of fermionic DQME-SQ
For fermionic environments, the DQME-SQ theory can be established in a similar way. For a single-level system, we consider a common bilinear form of system-environment coupling, represented by $H_{\mbox{\tiny S}\mbox{\tiny E}}=\hat{c}^{+}\hat{F}^{-}+\hat{F}^{+}\hat{c}^{-}$, which describes the exchange of fermionic particles between the system and environment. Here, $\hat{c}^{-}\equiv\hat{c}$ ($\hat{c}^{+}\equiv\hat{c}^{\dagger}$) denotes the electron annihilation (creation) operator in the system state. The DD for the environment hybridization operators reads
@@
Here, $\langle\hat{f}_{k}^{\sigma}(t)\hat{f}_{k^{\prime}}^{\bar{\sigma}}(0)\rangle_{\mbox{\tiny D}}=\delta_{kk^{\prime}}\eta_{k}^{\sigma}e^{-\gamma_{k}^{\sigma}t}$ with $\sigma=\pm$ and $\bar{\sigma}=-\sigma$. Such a DD protocol completely recovers the hybridization correlation functions. The fermionic dissipaton Fock space is ${\cal H}_{\mbox{\tiny D}}=\otimes_{k}({\cal H}_{f^{-}_{k}}\otimes{\cal H}_{f^{+}_{k}})$, which involves both left and right actions of dissipaton creation and annihilation operators. Thus, the second quantization of fermionic dissipatons is realized through the following mappings.
@@
where $\hat{N}_{k}\equiv\hat{b}_{k}^{+}\hat{b}_{k}^{-}$ is the fermionic dissipaton number operator. The detailed derivation of Eq. (III.1) and its formal equivalence to the fermionic HEOM are presented in Sec. III.2. Similar to the bosonic DQME-SQ, Eq. (III.1) can be straightforwardly extended to more complex OQSs; see Sec. III.3.
III.2 Exactness of fermionic DQME-SQ
Similar to the bosonic case, to verify the exactness of the fermionic DQME-SQ theory, we shall give a detailed derivation of Eq. (III.1) by starting from the fermionic HEOM formalism, which is known to be exact. For the quantum impurity model described above, the fermionic HEOM reads Jin08234703 ; Yan16110306
$\dot{\rho}_{{\bm{j}}}^{(n)}$ $=-\bigg{(}i{\cal L}_{\mbox{\tiny S}}+\sum_{r=1}^{n}\gamma_{j_{r}}\bigg{)}\rho_{{\bm{j}}}^{(n)}-i\sum_{\ell}{\cal A}^{\bar{\sigma}}\rho_{{\bm{j}}\ell}^{(n+1)}$
$\quad-i\sum_{r=1}^{n}(-)^{n-r}{\cal C}_{j_{r}}\rho_{{\bm{j}}_{r}^{-}}^{(n-1)}.$ (16)
Here, ${\cal L}_{\mbox{\tiny S}}\cdot\equiv[H_{\mbox{\tiny S}},\,\cdot\,]$, $j\equiv(k,\sigma)$, and ${\cal A}^{\bar{\sigma}}$ and ${\cal C}_{j}\equiv{\cal C}_{k}^{\sigma}$ are Grassmann superoperators defined by
${\cal A}^{\sigma}\hat{O}^{(n)}\equiv\hat{c}^{\sigma}\hat{O}^{(n)}+(-)^{n}\hat{O}^{(n)}\hat{c}^{\sigma}$ (17)
and
${\cal C}^{\sigma}_{k}\hat{O}^{(n)}\equiv\eta_{k}^{\sigma}\hat{c}^{\sigma}\hat{O}^{(n)}-(-)^{n}\eta_{k}^{\bar{\sigma}\ast}\hat{O}^{(n)}\hat{c}^{\sigma}.$ (18)
Here, $\hat{O}^{(n)}$ denotes $\rho^{(2m)}$ or $\rho^{(2m+1)}$ with even or odd fermionic parity, respectively. In the occupation number representation, we split the indices $\sigma=+$ and --, and rewrite the reduced density operator and auxiliary density operators as
$\rho_{\bm{j}}^{(n)}\longmapsto\rho_{{\bf n}{\bf m}}\equiv\rho_{n_{1}n_{2}\cdots n_{K};m_{1}m_{2}\cdots m_{K}},$ (19)
where $n_{k}=1$ if $(k,+)\in{\bm{j}}$; and otherwise $n_{k}=0$. Similarly, $m_{k}=1$ when $(k,-)\in{\bm{j}}$; and otherwise $m_{k}=0$. Consequently, the HEOM of Eq. (III.2) can be recast as
$\dot{\rho}_{{\bf n}{\bf m}}=-\big{(}i{\cal L}_{\mbox{\tiny S}}+\gamma_{{\bf n}{\bf m}}\big{)}\rho_{{\bf n}{\bf m}}$
$\quad-i\sum_{k}\Big{[}(-)^{M+N-\theta^{+}_{k}}\hat{c}^{-}\rho^{+-}_{{\bf n}_{k}^{+}{\bf m}}-(-)^{\theta^{+}_{k}}\rho^{+-}_{{\bf n}_{k}^{+}{\bf m}}\hat{c}^{-}$
$\quad+(-)^{M-\theta^{-}_{k}}\hat{c}^{+}\rho_{{\bf n}{\bf m}_{k}^{+}}-(-)^{N+\theta^{-}_{k}}\rho_{{\bf n}{\bf m}_{k}^{+}}\hat{c}^{+}$
$\quad+(-)^{M+N-\theta^{+}_{k}}\eta_{k}^{+}\hat{c}^{+}\rho_{{\bf n}_{k}^{-}{\bf m}}-(-)^{\theta^{+}_{k}-1}(\eta_{k}^{-})^{\ast}\rho_{{\bf n}_{k}^{-}{\bf m}}\hat{c}^{+}$
$\quad+(-)^{M-\theta^{-}_{k}}\eta_{k}^{-}\hat{c}^{-}\rho_{{\bf n}{\bf m}_{k}^{-}}-(-)^{N-1+\theta^{-}_{k}}(\eta_{k}^{+})^{\ast}\rho_{{\bf n}{\bf m}_{k}^{-}}\hat{c}^{-}\Big{]}.$ (20)
Here, we denote $\gamma_{{\bf n}{\bf m}}\equiv\sum_{k=1}(n_{k}\gamma_{k}^{+}+m_{k}\gamma_{k}^{-})$, $\theta^{+}_{k}=\sum_{l=1}^{k}n_{l}$ and $\theta^{-}_{k}=\sum_{l=1}^{k}m_{l}$, with $N=\theta_{K}^{+}$ and $M=\theta_{K}^{-}$. In deriving Eq. (III.2), we have used the relations
$\rho_{\bm{j};(k,+)}^{(n+1)}$ $\longmapsto(-)^{\theta^{-}_{K}+\theta^{+}_{K}-\theta^{+}_{k}}\rho_{n_{1}n_{2}\cdots n_{k}+1\cdots n_{K};m_{1}m_{2}\cdots m_{K}}$
$\quad\quad=(-)^{M+N-\theta^{+}_{k}}\rho_{{\bf n}_{k}^{+}{\bf m}},$
and
$\rho_{\bm{j};(k,-)}^{(n+1)}$ $\longmapsto(-)^{\theta^{-}_{K}-\theta^{-}_{k}}\rho_{n_{1}n_{2}\cdots n_{K};m_{1}m_{2}\cdots m_{k}+1\cdots m_{K}}$
$\quad\quad=(-)^{M-\theta^{-}_{k}}\rho_{{\bf n}{\bf m}_{k}^{+}}.$
To obtain Eq. (III.1), we introduce the fermionic dissipaton creation and annihilation operators, $\{\hat{b}^{\pm}_{k}\}$, which satisfy the standard fermionic anti-commutation relations. They also anti-commute with $\{\hat{c}^{\pm}\}$. The RDT $\tilde{\rho}(t)$ is defined as
$\tilde{\rho}(t)=\sum_{\bf mn}\bar{\rho}_{\bf nm}(t)|{\bf m}{\bf n}\rangle\rangle.$ (21)
Here,
$\bar{\rho}_{\bf nm}(t)=\prod_{kj}(\zeta_{k}^{-})^{-m_{k}}(\zeta_{j}^{+})^{-n_{j}}\rho_{\bf nm}(t),$ (22)
where $\rho_{\bf nm}(t)$ follow Eq. (III.2), $\zeta_{k}^{\sigma}=(\eta_{k}^{\sigma}\eta_{k}^{\bar{\sigma}\ast})^{\frac{1}{4}}$ with $\sigma=\pm$, $|{\bf mn}\rangle\rangle=|{\bf m}\rangle\langle{\bf n}|$, and
$|{\bf m}\rangle\equiv(\hat{b}^{+}_{1})^{m_{1}}\cdots(\hat{b}^{+}_{K})^{m_{K}}|{\bf 0}\rangle$ (23)
$|{\bf n}\rangle\equiv(\hat{b}^{+}_{K})^{n_{K}}\cdots(\hat{b}^{+}_{1})^{n_{1}}|{\bf 0}\rangle.$ (24)
By substituting Eqs. (III.2) and (22) into Eq. (21), we arrive at the fermionic DQME-SQ of Eq. (III.1).
III.3 Universality and compactness
It is straightforward to extend Eq. (III.1) to general OQSs containing multiple degrees of freedom that are coupled to more than one environment, without any loss of exactness. Consider the scenario of $H_{\mbox{\tiny S}\mbox{\tiny E}}=\sum_{u}\hat{c}_{u}^{+}\hat{F}_{\alpha u}^{-}+\hat{F}_{\alpha u}^{+}\hat{c}_{u}^{-}$, where $\alpha$ and $u$ label the environments and the system’s degrees of freedom, respectively. The fermionic DQME-SQ reads
$\dot{\tilde{\rho}}$ $=-i[H_{\mbox{\tiny S}},\tilde{\rho}]-\sum_{\alpha uk}\big{(}\gamma^{-}_{\alpha uk}\hat{N}_{\alpha uk}\tilde{\rho}+\gamma^{+}_{\alpha uk}\tilde{\rho}\hat{N}_{\alpha uk}\big{)}$
$\quad-i\sum_{\alpha uk}\!\Big{[}\zeta_{\alpha uk}^{-}\big{(}\hat{c}_{u}^{+}\hat{b}^{-}_{\alpha uk}\tilde{\rho}-\!\hat{b}_{\alpha uk}^{-}\tilde{\rho}\hat{c}_{u}^{+})$
$\quad\quad\quad\ +\zeta_{\alpha uk}^{+}(\hat{c}_{u}^{-}\tilde{\rho}\hat{b}^{+}_{\alpha uk}\!-\!\tilde{\rho}\hat{b}^{+}_{\alpha uk}\hat{c}_{u}^{-}\!\big{)}$
$\quad\quad\quad\ +\xi^{+}_{\alpha uk}\hat{c}_{u}^{+}\tilde{\rho}\hat{b}^{-}_{\alpha uk}-\xi^{+*}_{\alpha uk}\hat{b}^{+}_{\alpha uk}\tilde{\rho}\hat{c}_{u}^{-}$
$\quad\quad\quad\ -\xi^{-}_{\alpha uk}\hat{c}_{u}^{-}\hat{b}^{+}_{\alpha uk}\tilde{\rho}+\xi^{-*}_{\alpha uk}\tilde{\rho}\hat{b}^{-}_{\alpha uk}\hat{c}_{u}^{+}\Big{]},$ (25)
where the coefficients are chosen as $\zeta_{\alpha uk}^{\sigma}=(\eta_{\alpha uk}^{\sigma}\eta_{\alpha uk}^{\bar{\sigma}\ast})^{\frac{1}{4}}$ and $\xi_{\alpha uk}^{\sigma}=\eta_{\alpha uk}^{\sigma}/\zeta_{\alpha uk}^{\sigma}$.
The pseudofermion theory Cir23033011 presents a fermionic variant of the pseudomode approach. This theory utilizes a Hilbert space with a dimension identical to that associated with Eq. (III.1), given by $\otimes_{k}({\cal H}_{f^{-}_{k}}\otimes{\cal H}_{f^{+}_{k}})$. While it is speculated that a connection between the pseudofermion and DQME-SQ theories exists, directly establishing this relationship is challenging. The difficulty arises from the fact that the pseudofermion Lindblad equation involves projection operators associated with the even and odd fermionic parity spaces, as shown in Eqs. (16)-(17) of Ref. Cir23033011 . In contrast, this parity-based structure is not required by the DQME-SQ theory.
The compactness of the fermionic DQME-SQ arises from the direct construction of fermionic dissipatons based on the exponential decomposition of hybridization correlation functions, which generally results in complex-valued coefficients and exponents. On the contrary, the pseudofermion theory proposed in Ref. Cir23033011 requires an additional post-processing procedure to accommodate the complex-valued coefficients from the exponential decomposition. The post-processing procedure typically increases the number of pseudomodes required, which inevitably makes the numerical simulation more expensive. Such a post-processing procedure is not needed in the implementation of the DQME-SQ approach.
The fermionic DQME-SQ approach can be directly applied to investigate properties related to the environment hybridization operators, ${\hat{F}_{u}^{\pm}}$. For example, for the electric current defined as $\hat{I}=-i\sum_{u}\big{(}\hat{c}_{u}^{+}\hat{F}_{u}^{-}-\hat{F}_{u}^{+}\hat{c}_{u}^{-}\big{)}$, its expectation value is evaluated using the DQME-SQ theory as
$I(t)=i\sum_{uk}{\rm tr_{\mbox{\tiny S}}}\big{[}c_{u}^{-}\langle{\bm{0}}|\tilde{\rho}(t)|{\bm{0}}^{+}_{uk}\rangle-\langle{\bm{0}}^{+}_{uk}|\tilde{\rho}(t)|{\bm{0}}\rangle c_{u}^{+}\big{]}.$ (26)
@@ III.3 Universality and compactness
IV Representing DQME-SQ by quantum circuits
The DQME-SQ offers a notable advantage: its formulation embraces a time-local structure within a compact Hilbert space. This unique feature enables easy encoding into quantum circuits, distinguishing it from HEOM. While the computational complexity remains comparable to HEOM when implemented with classical algorithms, its excellent compatibility with quantum algorithms makes DQME-SQ a preferred choice for simulating complex non-Markovian OQSs.
To simulate the non-Markovian open quantum dynamics governed by the DQME-SQ using quantum circuits, the first key step is encoding the RDT. For bosonic environments, the RDT $\tilde{\rho}$ is mapped onto a pure qubit state, $\lvert\tilde{\rho}\rangle$, as
$\tilde{\rho}\longmapsto|\tilde{\rho}\rangle=\frac{1}{Z}\sum_{ij{\bf n}}P_{ij}^{\bf n}\,|i\rangle\otimes|j\rangle\otimes|\bf n\rangle.$ (27)
Here, the summation is over all the possible configurations of the dissipaton-embedded system, where $|{\bf n}\rangle=|n_{1}\rangle\otimes|n_{2}\rangle\cdots\otimes|n_{K}\rangle$ represents a dissipaton configuration. Note that for fermionic environments, $|{\bf n}\rangle$ is replaced by paired dissipaton Fock states [cf. Eq. (30)].
When implementing the DQME-SQ approach, it is necessary to truncate the dissipaton configurations. This truncation is particularly important in the case of bosonic environments, where the occupation number for a bosonic dissipaton level can, in principle, take on any value. In the context of digital quantum simulations, the truncation of dissipaton configurations effectively reduces the dimension of the RDT, which in turn alleviates the computational burden associated with the DQME-SQ implementation. The truncation of dissipaton configurations in the DQME-SQ approach is, in essence, equivalent to the truncation of the HEOM formalism. The latter has been extensively studied in the literature, and a variety of truncation schemes have been developed Tan06082001 ; Zha15214112 ; Han18234108 ; Zha21905 ; Din22224107 ; Zha23014106 . These established truncation schemes typically allow for rapid convergence of the results. However, in practice, it is still desirable to conduct careful convergence tests to ensure that the truncation errors are negligibly small.
The dissipaton configurations are encoded onto the qubits using a standard binary code. The reduced density operator of the original system under investigation is given by $\rho_{\mbox{\tiny S}}(t)=\langle\bm{0}|\tilde{\rho}(t)$, and its trace is preserved during the propagation of the DQME-SQ. For the representation of the RDT on qubits, a renormalization procedure is required due to the normalization of the qubits’ state. The global normalization coefficient $Z$ is determined by the probability conservation condition $\sum_{i}P_{ii}^{\bf 0}=1$. It is important to note that this renormalization procedure does not cause additional computational burden or introduce numerical errors during the quantum simulation; see Appendix B for further discussions.
The temporal evolution of $\tilde{\rho}(t)$ is then simulated by propagating the following equation of motion,
@@ IV Representing DQME-SQ by quantum circuits
where $\Lambda$ is a non-Hermitian dynamical generator that represents the right-hand side of Eq. (II.1) or Eq. (III.1). It reads
$\Lambda$ $=({H}_{\mbox{\tiny S}}\otimes I_{\mbox{\tiny S}}-I_{\mbox{\tiny S}}\otimes H_{\mbox{\tiny S}}^{\rm T})\otimes I_{\mbox{\tiny D}}\!-iI_{\mbox{\tiny S}}\otimes I_{\mbox{\tiny S}}\otimes\sum_{k}\!\gamma_{k}\hat{b}_{k}^{+}\hat{b}_{k}^{-}$
$\quad\!+\!({\hat{Q}}\otimes I_{\mbox{\tiny S}}-I_{\mbox{\tiny S}}\otimes\hat{Q}^{\rm T})\otimes\sum_{k}\zeta_{k}(\hat{b}_{k}^{-}+\hat{b}_{k}^{+})$
$\quad+i({\hat{Q}}\otimes I_{\mbox{\tiny S}}+I_{\mbox{\tiny S}}\otimes\hat{Q}^{\rm T})\otimes\sum_{k}\xi_{k}\hat{b}_{k}^{+}.$ (29)
Here, $I_{\mbox{\tiny S}}$ is the identity operator of the system Hilbert space ${\cal H}_{\mbox{\tiny S}}$, $I_{\mbox{\tiny D}}$ is identity operator of the dissipaton Hilbert space ${\cal H}_{\mbox{\tiny D}}$, and $\hat{O}^{\rm T}$ denotes the transpose of any operator $\hat{O}$. We then represent the state vector $|\tilde{\rho}\rangle$ on the qubits by using the standard binary code. Equations (27) and (IV) constitute the encoding protocol of the bosonic DQME-SQ.
The fermionic counterpart of Eq. (27) reads
$\tilde{\rho}\longmapsto|\tilde{\rho}\rangle=\frac{1}{Z}\sum_{\mu\nu{\bf mn}}P_{\mu\nu}^{\bf mn}|\mu\rangle\otimes|\nu\rangle\otimes|{{\bf m}}\rangle\otimes|{\bf n}\rangle.$ (30)
Here, $|\mu\rangle,|\nu\rangle\in\mathcal{H}_{\mbox{\tiny S}}$ are the system states. The dynamical generator of Eq. (III.1) can be similarly constructed as
$\Lambda=({H}_{\mbox{\tiny S}}\otimes I_{\mbox{\tiny S}}-I_{\mbox{\tiny S}}\otimes H_{\mbox{\tiny S}}^{\rm T})\otimes I_{\mbox{\tiny D}}\otimes I_{\mbox{\tiny D}}$
$-iI_{\mbox{\tiny S}}\otimes I_{\mbox{\tiny S}}\otimes\Big{(}\sum_{k}\gamma^{-}_{k}\hat{N}_{k}\otimes{I}_{\mbox{\tiny D}}+{I}_{\mbox{\tiny D}}\otimes\sum_{k}\gamma^{+}_{k}\hat{N}_{k}\Big{)}$
$+\sum_{k}\!\Big{[}\zeta^{-}_{k}\big{(}\hat{c}^{+}\otimes I_{\mbox{\tiny S}}\otimes\hat{b}_{k}^{-}\otimes{I}_{\mbox{\tiny D}}-I_{\mbox{\tiny S}}\otimes\hat{c}^{-}\otimes\hat{b}_{k}^{-}\otimes{I}_{\mbox{\tiny D}}\big{)}$
$+\zeta^{+}_{k}\big{(}\hat{c}^{-}\!\otimes I_{\mbox{\tiny S}}\otimes{I}_{\mbox{\tiny D}}\otimes\hat{b}_{k}^{-}-I_{\mbox{\tiny S}}\otimes\hat{c}^{+}\!\otimes{I}_{\mbox{\tiny D}}\otimes\hat{b}_{k}^{-}\big{)}$
$+\xi^{+}_{k}\hat{c}^{+}\otimes I_{\mbox{\tiny S}}\otimes{I}_{\mbox{\tiny D}}\otimes\hat{b}_{k}^{+}-\xi^{+\ast}_{k}I_{\mbox{\tiny S}}\otimes\hat{c}^{+}\otimes\hat{b}_{k}^{+}\otimes{I}_{\mbox{\tiny D}}$
$-\xi^{-}_{k}\hat{c}^{-}\otimes I_{\mbox{\tiny S}}\otimes\hat{b}_{k}^{+}\otimes{I}_{\mbox{\tiny D}}+\xi^{-\ast}_{k}I_{\mbox{\tiny S}}\otimes\hat{c}^{-}\otimes{I}_{\mbox{\tiny D}}\otimes\hat{b}_{k}^{+}\Big{]}.$ (31)
The two parts, $|\mu\rangle\otimes|\bf{m}\rangle$ and $|\nu\rangle\otimes|{{\bf n}}\rangle$ in $|\tilde{\rho}\rangle$ of Eq. (30), are then encoded separately with different sets of Jordan-Wigner codes. Equations (30) and (IV) constitute the encoding protocol of the fermionic DQME-SQ.
The temporal propagation of $|\tilde{\rho}(t)\rangle$ can be realized with various quantum algorithms, including the linear combination of unitaries (LCU) Lon06825 ; Sch22023216 , quantum imaginary time evolution Kam22010320 ; Mot20205 , quantum singular value transformation Low17010501 ; Low19163 ; Mar21040203 ; Sur231002 , time-dependent variational principle Sch21270503 ; Sch22023216 ; Hea21013182 ; Wan234851 , and others Lam23Arxiv2310_12539 ; Zha24054101 ; Sur231002 . Here, we utilize the LCU propagation Sch22023216 for the demonstration. Within a small time step $\delta t$, the propagator $e^{-i{\Lambda}\delta t}$ is approximately decomposed as $e^{-i{\Lambda}\delta t}\approx U_{1}(\delta t)U_{0}(\delta t)$. Here, $U_{0}(\delta t)\equiv e^{-i\Lambda_{0}\delta t}$ governed by a Hermitian generator, $\Lambda_{0}\equiv(\Lambda+\Lambda^{\dagger})/2$, is a unitary propagator, while $U_{1}(\delta t)\equiv e^{-i\Lambda_{1}\delta t}$ governed by an anti-Hermitian generator, $\Lambda_{1}\equiv(\Lambda-\Lambda^{\dagger})/2$, is non-unitary. The latter is further approximated by Sch21270503
@@
Shown in Fig. 2 is the designed quantum circuit of DQME-SQ comprising the target bits to store the RDT state as well as a control bit for propagation. The input of the circuit is $|\tilde{\rho}(t)\rangle\otimes|0\rangle$. From the left to the right in the circuit, we first act a Hadamard gate (H) on the control bit, followed by the $U_{0}(\delta t)$ gate on the target bits. Subsequently, the states go through the 0-control and 1-control gates of $U^{\epsilon}_{\pm}(\delta t)$ successively. After acting an ${\rm R}_{\rm Y}(-{\pi}/{2})$ gate on the control bit, we obtain $2\epsilon|\tilde{\rho}(t+\delta t)\rangle\otimes|0\rangle+|{\rm unwanted}\rangle\otimes|1\rangle$. Finally, by performing an Z-measurement on the control bit, we project out the unwanted state and obtain the target state $|\tilde{\rho}(t+\delta t)\rangle$ with a probability depending on $\epsilon$. Once obtained, the wanted state will serve as the initial state for the next propagation step.
@@
By using the Trotter decomposition, the unitary operators $U_{0}(\delta t)$ and $U_{\pm}^{\epsilon}(\delta t)$ can be split into a series of simpler quantum logic units. This facilitates the practical implementation of the DQME-SQ theory. In contrast to classical algorithms, whose computational costs grow exponentially as the system size or the complexity of non-Markovianity increases Tan20020901 , the LCU-Trotter algorithm exhibits highly favorable scalability. Specifically, both the number of qubits and the circuit depth grow linearly with the increase in the degrees of freedom of the dissipatons. This favorable scalability makes the DQME-SQ theory, implemented with the LCU-Trotter algorithm, as a promising approach to efficiently simulate complex non-Markovian open quantum dynamics.
It should be mentioned that the unitary decomposition approximation in Eq. (32) will incur a systematic error on the order of $O(\epsilon^{2})$ Sch22023216 . For practical applications on real quantum computers in the noisy intermediate-scale quantum (NISQ) regimes and beyond, these errors and measurement costs from the unitary decomposition approximations can be significant Sch21270503 ; Sch22023216 . A sufficiently large $\epsilon$ is needed to resolve the matrices representing the quantum gates from the device noise, while a small $\epsilon$ is needed to obtain an accurate expansion of Eq. (32). To address this conflict, techniques such as Richardson’s extrapolation can be employed to improve the epsilon error and generate results with higher-order accuracy Sch21270503 ; Sch22023216 . Additionally, amplitude amplification technology can be used to reduce the measurement costs by amplifying the probability of detecting the target state Bra0253 ; Kwo21062438 ; Sur231002 . Therefore, the errors and measurement costs induced by the unitary decomposition approximation in Eq. (32) are thus controllable and expected to be gradually alleviated as quantum computing technologies advance.
@@
Figure 3: (a) Simulated time evolution of $P_{1}-P_{0}$, the population difference between two states, in a spin-boson model, by using the quantum algorithm for the DQME-SQ theory, compared to low and high temperature results via the classical propagation of HEOM; (b) Simulated time evolution of the ground-state population $P_{0}$ for a single-impurity Anderson model, with inset depicting the imaginary part of the correlation function. Detailed parameters are provided in the Appendix C.
V Numerical demonstration through digital quantum simulation